OpenStudy (anonymous):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 3x - 2. { -8, -2, 4 } { 8, -2, 4 } { 12, 0, -12 } none of the above

OpenStudy (anonymous):

its just put and get put the domain values in y and get range

OpenStudy (anonymous):

I got C

OpenStudy (anonymous):

do u want solution

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ist case y=3x-2 =3(-2)-2=-8 2nd y=3x-2 =3(0)-2 =-2 3rd y=3x-2 =3(2)-2=4

OpenStudy (anonymous):

love u